The claim is that the proportion of peas with yellow pods is equal to 0.25​ (or 25%). The sample statistics from one experiment include 570570 peas with 169169 of them having yellow pods. Find the value of the test statistic.
Answer by Boreal(15235)

The test stat is (phat-0.25)/sqrt{p*(1-p)/n}=(0.296-0.250)/0.0181=2.54